Section 12

Probability and Statistics

In case of several authors, Invited Speakers are marked with a ^{∗}.

David J. Aldous: Stochastic Coalescence . . . . . . . . . . . . . . . . . . . . . . . . . . . III 205
Maury Bramson: State Space Collapse for Queueing Networks . . . . . III 213

Mark I. Freidlin: Random and Deterministic Perturbations of Nonlinear Oscillators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 223

Jayanta K. Ghosh: Bayesian Density Estimation . . . . . . . . . . . . . . . . . . . III 237

F. Götze: Lattice Point Problems and the Central Limit Theorem in Euclidean Spaces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 245

Peter Hall^{∗} and Brett Presnell: Applications of Intentionally Biased Bootstrap Methods . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 257

Iain M. Johnstone: Oracle Inequalities and Nonparametric Function Estimation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 267

Jean-François Le Gall: Branching Processes, Random Trees and Superprocesses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 279

David Siegmund: Genetic Linkage Analysis: an Irregular Statistical Problem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 291

Alain-Sol Sznitman: Brownian Motion and Random Obstacles …. III 301

Boris Tsirelson: Within and Beyond the Reach of Brownian Innovation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 311

R. J. Williams: Reflecting Diffusions and Queueing Networks . . . . . . . III 321

Documenta Mathematica · Extra Volume ICM 1998 · III ·

204

Documenta Mathematica · Extra Volume ICM 1998 · III ·